Ammonia has been studied as an alternative "clean" fuel for internal combustion engines, since its reaction with oxygen produces only nitrogen and water vapor, and in the liquid form it is easily transported. An industrial chemist studying this reaction fills a 5.0 L flask with 4.7 atm of ammonia gas and 2.7 atm of oxygen gas at 48.0 Celsius. He then raises the temperature, and when the mixture has come to equilibrium measures the partial pressure of nitrogen gas to be 1.6 atm. Calculate the pressure equilibrium constant for the combustion of ammonia at the final temperature of the mixture. Round your answer to significant digits.

K_p=?
Answer #1

4 NH3 aN GO initial a 3x at eqili:4.7- 4x ax 1 O.8 ath 6 (6x) (ax) Kp= (N(Mo X 4 3 (4.7-4x(27-3x) X ) 31310 3158 Kp= 2.4906 x